Three years ago, Max's age was one-fourth Betty's age (b). Write an expression to show max' safe today.
Answer by stanbon(75887)

Three years ago, Max's age was one-fourth Betty's age (b). Write an expression to show max' age today.
-----
Equations:
m-3 = (1/4)b
----
Ans: m = (1/4)b + 3
------------------------
Cheers,
Stan H.
------------------